you have 12 bills in your wallet some of the bills are $5's and some are $10's totaling $100 how many of ech do you have? I am not very good with figuring out word problems can you help?
Answer by palanisamy(496)

Let the number of $5 bills = x
And the number of $10 bills = y
Total nomber of bills x+y = 12 ...(1)
Total value is 5x+10y = 100 ...(2)
(1)*5=> 5x+5y = 60 ...(3)
(2)-(3)=> 5y=40
y= 40/5
y= 8
(1)=> x+8=12
x=12-8
x=4
So, the number of $5 bills = 4
And the number of $10 bills = 8